OpenStudy (anonymous):

There are 8 crackers in 1 serving, 16 crackers in 2 servings, 24 crackers in 3 servings, and so on. How many crackers are in a box of 12 servings? A.80 B.96 C.108 D.120

OpenStudy (anonymous):

96 bc 12 times 8 is 96

OpenStudy (anonymous):

so B

OpenStudy (texaschic101):

8/1 = x/12 --- 8 crackers to 1 serv = x crackers to 12 serv cross multiply 1x = (8)(12) x = 96 I agree...96 crackers for 12 servings

OpenStudy (micheal1234):

thanks